Sandrine Machinery is a Swiss multinational manufacturing company. Currently, Sandrine's financial planners are considering undertaking a 1-year project in the United States. The project's expected dollar-denominated cash flows consist of an initial investment of $2000 and a cash inflow the following year of $2400. Sandrine estimates that its risk-adjusted cost of capital is 14%. Currently, 1 U.S. dollar will buy 0.78 Swiss franc. In addition, 1-year risk-free securities in the United States are yielding 6.75%, while similar securities in Switzerland are yielding 3%. Do not round intermediate calculations.

  1. If this project was instead undertaken by a similar U.S.-based company with the same risk-adjusted cost of capital, what would be the net present value and rate of return generated by this project? Round your answers to two decimal places.

    NPV = $  

    Rate of return = %

  2. What is the expected forward exchange rate 1 year from now? Round your answer to two decimal places.

    SF per U.S. $

  3. If Sandrine undertakes the project, what is the net present value and rate of return of the project for Sandrine? Do not round intermediate calculations. Round your answers to two decimal places.

    NPV = Swiss Francs

    Rate of return = %
Answer #1

a). NPV = PV of Cash Inflows - PV of Cash Outflows

= [$2,400 / (1 + 0.14)] - $2,000 = $2,105.26 - $2,000 = $105.26

Rate of Return = [Cash Flow in 1 year / Initial Investment] - 1

= [$2,400 / $2,000] - 1 = 1.20 - 1 = 0.20, or 20%

b). Forward Rate = Spot Rate * [{1 + r(Swiss)} / {1 + r(USD)}]

= 0.78 * [(1 + 0.03) / (1 + 0.0675)]

= 0.78 * 0.9649 = SF 0.75 / $

c). Cash flow year 1(Swiss Franc) = Cash Flow(USD) * Forward Rate

= $2,400 * 0.75 = SF 1,806.24

Initial Investment(Swiss Franc) = Initial Investment(USD) * Spot Rate

= $2,000 * 0.78 = SF 1,560

NPV = PV of Cash Inflows - PV of Cash Outflows

= [SF 1,806.24 / (1 + 0.14)] - $2,000 = SF 1,584.42 - SF 1,560 = SF 24.42

Rate of Return = [Cash Flow in 1 year / Initial Investment] - 1

= [SF 1,806.24 / SF 1,560] - 1 = 1.1578 - 1 = 0.1578, or 15.78%